Two forces F1 and F2 act on a particle. F1 has magnitude 6N and in direction 45°, and F2 has magnitude 7N and in direction 154°. Find the magnitude and direction of their resultant

1
Expert's answer
2021-12-27T13:02:37-0500

The resultant force

"{\bf F}={\bf F}_1+{\bf F}_2"

In projections

"F_x=F_{1x}+F_{2x}=F_1\cos \theta_1+F_2\cos\theta_2\\
F_y=F_{1y}+F_{2y}=F_1\sin \theta_1+F_2\sin\theta_2"

"F_x=6\cos 45^\circ+7\cos154^\circ=-2.0\\
F_y=6\sin 45^\circ+7\sin154^\circ=7.3"

The magnitude of resultant force

"F=\sqrt{F_x^2+F_y^2}=\sqrt{(-2.0)^2+7.3^2}=7.6"

The direction of resultant force

"\theta=\tan^{-1}\frac{F_y}{F_x}=\tan^{-1}\frac{7.3}{-2.0}=105^\circ"